OpenStudy (anonymous):

What would a 2x3 matrix look like? Example please.

OpenStudy (anonymous):

its row by column so \[\left[\begin{matrix}? & ? & ? \\ ? & ? & ? \\ \end{matrix}\right]\]

OpenStudy (amistre64):

2rows; 3cols .. yep

OpenStudy (amistre64):

\[\left[ \begin{array}c 1 & 2 & 3\\4 & 3 & 2\\\end{array} \right]\]